A metal alloy that is made up of 15% aluminum is combined with a metal alloy
that is 35% aluminum. If the result is 400 Kilograms of 27.5% of aluminum alloy, then
how much of each alloy was used?

Answer by ikleyn(53765)

.
Math model is this system of two equations

    x +     y = 400,
0.15x + 0.35y = 0.275*400

Solve it by any method you know.


Alternatively, you can use the model with one equation

0.15x + 0.35*(400-x) = 0.275*400.

Here x is the amount of the 15% aluminium alloy, y is the amount of the 35% aluminium alloy.

There is entire bunch of lessons covering various types of mixture problems
    - Mixture problems
    - More Mixture problems
    - Solving typical word problems on mixtures for solutions
    - Word problems on mixtures for antifreeze solutions
    - Word problems on mixtures for dry substances like coffee beans, nuts, cashew and peanuts
    - Word problems on mixtures for dry substances like candies, dried fruits
    - Word problems on mixtures for dry substances like soil and sand
    - Word problems on mixtures for alloys (*)
    - Typical word problems on mixtures from the archive
    - OVERVIEW of lessons on word problems for mixtures
in this site. One of them, related to alloys, is marked by (*).

Read them and become an expert in solution the mixture word problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

It contains many other solved word problems, as well as many other interesting and useful things.